OpenStudy (anonymous):

ali while driving to school,computes the average speed for his trip to be 20 km/h.on his return trip along the same route there is less traffic and averege speed is 30 km/h.what is the average speed for ali's trip?

OpenStudy (anonymous):

the speed for going to is 20kmph and to return is 30kmph so the average speed =v1+v2/2 therefore v average=25kmph is his average speed

OpenStudy (kropot72):

Let the distance from home to school = d kilometers. Time to travel to school = d/20 hours Time to travel home from school = d/30 hours Total time for return trip = (d/20 + d/30) hours Total distance travelled for return trip = 2d kilometers \[Average\ speed=\frac{distance}{time}=\frac{2d}{\frac{d}{20}+\frac{d}{30}}=2d \times \frac{60}{5d}=\frac{120}{5}=24\ km/hr\]